OpenStudy (cmtboy2):

Davida bought a pair of earrings, each in the shape of an isosceles triangle with equal side lengths of 24 mm and a base length of 16 mm. She used one earring to make another pair of earrings, similar in shape, with the base length equal to 11 mm. What is the length of the equal-length sides of this new pair of earrings? mm

OpenStudy (cmtboy2):

@TheCuteOne25

OpenStudy (misssmartiez):

@retirEEd

OpenStudy (thecuteone25):

Multiply the sides together

OpenStudy (misssmartiez):

ohh i see what she is doing, yes try to multiply them, i was adding up the angles instead of *,

OpenStudy (misssmartiez):

tell me if you don't get the answer.

OpenStudy (cmtboy2):

4,224

OpenStudy (cmtboy2):

11x24x16=4,224

OpenStudy (thecuteone25):

ok never mind don't multiply

OpenStudy (cmtboy2):

ugh i swear you are hurd to do math with lol

OpenStudy (thecuteone25):

lol

OpenStudy (misssmartiez):

LOL

OpenStudy (misssmartiez):

http://brainly.com/question/352837 "If the base is 11 mm the side lengths are 19 mm "

OpenStudy (cmtboy2):

cheater

OpenStudy (thecuteone25):

yeah go to that sight dom cause I have no idea

OpenStudy (misssmartiez):

"Id say 19mm since the difference in length in the similarly shape earnings side length is 8" http://brainly.com/question/339726 Source, so I was correct, 11 as the base and 2 sides as 19 or 10.5, which I thought and it was 19 c:

OpenStudy (cmtboy2):

a right

OpenStudy (cmtboy2):

a or no

OpenStudy (thecuteone25):

I think it's A

OpenStudy (cmtboy2):

@MissSmartiez

OpenStudy (misssmartiez):

this can help watch video

OpenStudy (cmtboy2):

i can't

OpenStudy (cmtboy2):

@shannonhall

OpenStudy (misssmartiez):

oh well, here is an example problem. Solution: Step 1: If we multiply the length of one side of the larger rectangle by the scale factor we get the length of the corresponding side of the smaller rectangle. Step 2: Dimension of larger rectangle × scale factor = dimension of smaller rectangle Step 3: 24 × scale factor = 20 [Substitute the values.] Step 4: Scale factor = 20/24 [Divide each side by 24.] Step 5: Scale factor =5/6= 5:6 [Simplify.] Therefore, scale factor from the larger rectangle to the smaller rectangle is 5:6.

OpenStudy (shannonhall):

bro to be honest idk im stuck on my own prob

OpenStudy (cmtboy2):

lol

OpenStudy (misssmartiez):

lol.

OpenStudy (cmtboy2):

ok then @Qwertty123

OpenStudy (misssmartiez):

Ok just do 21/.57 = 11.97, rounded, 12. so I would say to the first question probably A, but still finding out the rest brb

OpenStudy (misssmartiez):

Yes, A is correct. the other questions don't add up to 12.

OpenStudy (thecuteone25):

do you get it @cmtboy2

OpenStudy (cmtboy2):

no

OpenStudy (thecuteone25):

ugggh

OpenStudy (misssmartiez):

Here, you turn each fraction into a decimal, by doing the less number divided by the greater number, like 4 divided by 7. Then round. Then use the calculator to divide that answer by 21. = 11.97, round that and it equals 12. So it is A. c:

OpenStudy (cmtboy2):

what one is a

OpenStudy (misssmartiez):

the first answer

OpenStudy (cmtboy2):

whats the last one

OpenStudy (cmtboy2):

@Qwertty123 can you help me plz

OpenStudy (qwertty123):

Umm hold on

OpenStudy (cmtboy2):

ok

OpenStudy (qwertty123):

scale factor: 14÷20=?

OpenStudy (cmtboy2):

0.7

OpenStudy (qwertty123):

YEAH! :D

OpenStudy (cmtboy2):

thats the answer

OpenStudy (qwertty123):

Yeah!

OpenStudy (cmtboy2):

thank you and i will picky

OpenStudy (cmtboy2):

lol